# Fast LiDAR Data Generation with Rectified Flows

### v1

@@ Abstract
Building LiDAR generative models holds promise as powerful data priors for restoration, scene manipulation, and scalable simulation in autonomous mobile robots. In recent years, approaches using diffusion models have emerged, significantly improving training stability and generation quality. Despite the success of diffusion models, generating high-quality samples requires numerous iterations of running neural networks, and the increasing computational cost can pose a barrier to robotics applications. To address this challenge, this paper presents R2Flow, a fast and high-fidelity generative model for LiDAR data. Our method is based on rectified flows that learn straight trajectories, simulating data generation with much fewer sampling steps against diffusion models. We also propose a efficient Transformer-based model architecture for processing the image representation of LiDAR range and reflectance measurements. Our experiments on the unconditional generation of the KITTI-360 dataset demonstrate the effectiveness of our approach in terms of both efficiency and quality.
@@ I Introduction
LiDAR sensors provide accurate 3D point clouds of surrounding environments by performing an omnidirectional time-of-flight (ToF) ranging. The LiDAR point clouds play an important role for autonomous mobile robots to understand their surroundings both geometrically and semantically, through techniques such as SLAM, object detection, and semantic segmentation. However, the performance of these techniques can be degraded due to incompleteness in adverse weather conditions and point density gaps between different LiDAR sensors. Restoring the degraded point clouds requires a data prior that models complex real-world patterns.
Generative modeling of LiDAR data [1, 2, 3, 4, 5, 6, 7, 8] has been studied for the purposes, motivated by the significant progress in deep generative models [9]. Deep generative models aim to build neural networks that represent the probability density distribution underlying the given samples, e.g., large-scale datasets of LiDAR point clouds. So far, the deep generative models trained on LiDAR data have proven useful as prior knowledge in translation tasks such as sparse-to-dense completion [2, 3, 5, 4] and simulation-to-real (sim2real) domain transfer [3], which also boost the perception tasks such as semantic segmentation. In particular, among the generative model frameworks, diffusion models have led to substantial advancements in LiDAR domain [5, 4, 6, 8] due to their training stability and ability to generate high-quality samples.
Despite the success of diffusion models, generating high-quality LiDAR data incurs significant computational cost. In general, sampling in diffusion models is formulated as a stochastic differential equation (SDE) [10] that describe trajectories from the latent distribution to the data distribution. Accurately simulating these learned trajectories requires hundreds or even thousands of discretized steps, each involving the execution of a deep neural network. Fig. 1 illustrates the trade-offs between the number of sampling steps and the sample quality for recent methods on LiDAR data generation [6, 4]. As shown in Fig. 1, naively reducing the number of sampling steps degrades the quality of the generated LiDAR data. This limitation poses a challenge for robotics applications, where power efficiency and computational speed are critical constraints.
(a) LiDM [6] + APE (b) R2DM [4] (c) R2Flow (ours)
Refer to caption Refer to caption Refer to caption
Feature-space iteration Pixel-space iteration Pixel-space iteration
273M params 31M params 81M params
Refer to captionblurry Refer to caption Refer to caption
NFE = 200 NFE = 256 NFE = 256
(3.7 s) (3.7 s) (7.6 s)
@@
NFE = 2 NFE = 2 NFE = 2
(38 ms) (56 ms) (62 ms)
Figure 1: Comparison of range image-based LiDAR generative models. Each column shows a schematic diagram of an iterative generation process (top) and LiDAR point clouds (middle & bottom) generated with the different NFEs (number of function evaluations). In comparison to the baselines (a & b), our proposed R2Flow (c) can generate the consistent and high-quality samples even with the low NFE setting.
To this end, we propose R2Flow (Range–Reflectance Flow), a novel generative model for fast and realistic LiDAR data generation. As a framework for building generative models, we employ rectified flows [11, 12], a kind of conditional flow matching framework for training continuous normalizing flows [13, 14], which have demonstrated generating natural images such as human faces and common objects. Similar to diffusion models, the rectified flows represent the data generation as an iterative process using a deep neural network. Its most distinctive feature is that the process is formulated as deterministic straight trajectories, while the diffusion models employ stochastic curved trajectories. Therefore, the sampling is robust to the number of steps or the step size, and ideally, the entire generation trajectory can be simulated in a single step. Similar to relevant studies [1, 2, 3, 4, 5, 6], R2Flow is trained on the equirectangular image representation of multimodal measurements: range and reflectance (intensity of laser reflection). Moreover, we also propose a neural network architecture to generate the multimodal images based on lightweight vision transformers (ViTs) [15]. Among recent architectures, we verify that our approach achieve the better sample quality for the LiDAR data generation task, suppressing the computational cost and the model size. We conduct an unconditional generation task with the KITTI-360 dataset [16] to evaluate quality and efficiency. Our approach surpasses the state-of-the-art results for both large and small number of steps. We summarize our contributions as follows:
@@
  • We also propose a model architecture built upon the recent vision transformers, which is compatible of fidelity and efficiency for LiDAR data generation.
  • The effectiveness of our approach is demonstrated through an unconditional generation task on the KITTI-360 dataset.
@@ II Related Work
Here, we briefly summarize three classes of the existing LiDAR generative models using neural networks.
@@
Variational autoencoders (VAEs). VAEs [17] are trained with an autoencoder with latent representation at the bottleneck. Caccia et al. [1] initiated early work on generating LiDAR range images using a vanilla VAE. In spite of the ELBO-based stable training, the generated samples were blurry. More recently, Xiong et al. [7] employed the improved framework, vector-quantized variational autoencoder (VQ-VAE) [18], with the voxel representation.
@@
Generative adversarial networks (GANs). GANs [19] consist of two competing neural networks, a generator and a discriminator, and have been actively applied in various domains over the last decade. Caccia et al. [1] reported the first results by training a basic GAN on range images. DUSty [2] and DUSty v2 [3] proposed architectures robust to raydrop noise (missing points caused by non-returned laser signals). GANs achieved better quality than VAEs; however, training is unstable, and the resultant point clouds still deviate significantly from real samples.
@@
Diffusion models. In recent years, diffusion models have gained significant attention for their stable training and high-quality sample generation. Diffusion models define the data and latent spaces with the same dimensionality and construct a Markov chain between them over multiple timesteps. Moreover, it is known that the family of diffusion models can be generalized as stochastic differential equations (SDEs) [10]. Here we introduce some studies that proposed the diffusion models for LiDAR data generation. LiDARGen [5] trained range and reflectance images using a variance-exploding SDE [20] trained on pixel-space with a discrete-time schedule. R2DM [4] also trained range and reflectance images but utilized a variance-preserving SDE [21] trained on pixel-space with a continuous-time schedule. The pixel-space diffusion models can capture fine details while they require high computational cost due to the iterative nature of sampling. To mitigate this issue, Ran et al. [6] proposed LiDM with architectural improvements based on the latent diffusion model (LDM) [22]. LiDM first pre-trained an autoencoder to compress the range images and then trained a discrete-time diffusion model on the lower dimensional feature space. RangeLDM [8] is a more recent work following a similar LDM approach. Nevertheless, the LDM approaches still struggle with blurriness caused by the non-iterative decoding by the autoencoder.
@@
In summary, the diffusion model-based approach generates high-fidelity samples with stable training, among the generative model frameworks. However, the sampling process requires a sufficiently large number of steps because the generative process is defined by stochastic curved trajectories defined as SDEs. If these trajectories are approximated with too few steps, the generated LiDAR samples are prone to discretization errors (see Fig. 1 for an example). In this paper, we address this issue by introducing easy-to-approximate trajectories demonstrated in natural image domains [11].
@@ III Method
We employ rectified flow [11] and its extension [12] to construct the generative trajectories that are optimized to be straight so that sampling can be performed in few-step approximation. In this section, we first introduce the procedure to build the straight trajectories by rectified flows and then describe our modifications to LiDAR data generation.
@@ III-A Preliminary
Initial training. Suppose the unknown data distribution $p_{1}$ where the dataset samples $\bm{x}_{1}\sim p_{1}$ are only accessible and Gaussian distribution $p_{0}=\mathcal{N}(0,\bm{I})$ which draws the latent variables $\bm{x}_{0}\sim p_{0}$. Both $p_{1}$ and $p_{0}$ are defined over $\mathbb{R}^{d}$. The goal is to build the transport map between the two distributions. In rectified flows, we consider the following ordinary differential equation (ODE) which transports the samples within the timesteps $t\in[0,1]$:
@@
where $\bm{x}_{t}\in\mathbb{R}^{d}$ is an intermediate state at $t$ and $v_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a neural network to predict the velocity fields towards $\bm{x}_{1}$. To minimize the discretization errors in ODE integration, we consider the linear interpolation path between $\bm{x}_{0}$ and $\bm{x}_{1}$: $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$. Fig. 3(a) illustrates the example trajectory. Then we train the network $v_{\theta}$ by minimizing the following conditional flow matching loss $\mathcal{L}_{\rm{CFM}}$ for independently sampled pairs $(\bm{x}_{1},\bm{x}_{0})$, so that the network $v_{\theta}$ to force the sample $\bm{x}_{t}$ to follow the uniform velocity $\bm{x}_{1}-\bm{x}_{0}$ as much as possible [11].
@@
where $t\sim\mathrm{Uniform}(0,1)$ and $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$. The initial model $v_{\theta}$ obtained from Eq. 2 is referred to as 1-RF in this paper.
@@
Sampling. Sampling can be performed by solving the initial value problem of the ODE described in Eq. 1, using the learned velocity estimators $v_{\theta}$. For 1-RF and 2-RF, we can choose the arbitrary number of sampling steps. In general, the larger the number of steps, the smaller the discretization error. For $k$-TD models, the number of sampling steps is fixed at $k$. Any integration solvers can be used for sampling, such as the following Euler methods:
@@
Figure 2: Scene interpolation using R2Flow inversion. The both side were reconstructed from real samples via inversion. The middle four samples were generated from interpolated latent variables.
@@ III-B Data Representation
Following the existing studies [1, 2, 3, 4, 5, 6], R2Flow is trained on the equirectangular image representation of LiDAR data. We assume a LiDAR sensor that has an angular resolution of $W$ in azimuth and $H$ in elevation and measures the range and reflectance at each laser angle. Then, $HW$ sets of the range and reflectance values can be projected to a 2-channel equirectangular image $\bm{x}_{1}\in\mathbb{R}^{2\times{H}\times{W}}$ by spherical projection. Moreover, following the baselines [5, 4], we also re-scale the range modality $\bm{x}_{\rm{range}}\in[0,x_{\rm{max}}]^{1\times{H}\times{W}}$ into the log-scale representation $\bm{x}_{\rm{log}}\in[0,1]^{1\times{H}\times{W}}$ as follows:
@@ III-C Velocity Estimator
Pixel-space vs. feature-space. As discussed in Sec. II, the pixel-space iteration involves high computational cost in general, which barriers the adoption of powerful backbone models such as Vision Transformers (ViTs) [24]. A common approach to this issue is to reduce the dimensionality at which iterative models are trained. In diffusion models, Rombach et al. [22] proposed LDM (latent diffusion model), which consists of an autoencoder (AE) pretrained to perceptually compress images and a diffusion model trained on the lower-dimensional AE features. The compression is motivated by the observation that representing imperceptible details of natural images can be relegated to AE. However, the AE-based non-iterative decoding can be problematic in LiDAR range image generation which requires accurate pixel values and their alignment to maintain geometric fidelity in point clouds. For instance, LiDM [6] proposed the LDM-based approach in the LiDAR generation task but also identified blurry patterns output by the AE decoder as a remaining issue (see Figs. 1 and 6). Therefore, we rethink powerful yet efficient pixel-space architectures for precise modeling and aim to reduce the number of iteration as described in Section III-A.
Architecture design. Our model is built upon HDiT (hourglass diffusion transformer) [15], which is a ViT-based architecture proposed for pixel-space diffusion models. The key idea is to use a sliding window self-attention mechanism [23] to avoid increasing the $\mathcal{O}(n^{2})$ computation with respect to the number of tokens $n$. It thereby enables high-resolution image generation without using the aforementioned compression [22] or staged upsampling [25], despite the pure Transformer structure. In this paper, to generate panoramic range and reflectance images representing LiDAR data, we modify HDiT as follows. (i) The sliding window in the self-attention layers [23] is modified to operate in a horizontal circular pattern using a circular padding technique [5, 4]. (ii) Following the recent ViT-based architecture for LiDAR processing [26], the patch size in tokenization is changed from the default square shape to a landscape shape of $1\times 4$. The sliding windows also has the landscape shape of $3\times 9$. (iii) We use pre-defined LiDAR beam angles to condition the relative positional embeddings (RoPE [27]) in the self-attention layers, limiting the angular frequencies to harmonics. (iv) Similar to ViT [24], we apply a learnable additive bias to the tokens as an absolute positional embedding (APE); otherwise, the generated LiDAR point clouds involve random azimuth rotation.

### v2

@@ Abstract
Building LiDAR generative models holds promise as powerful data priors for restoration, scene manipulation, and scalable simulation in autonomous mobile robots. In recent years, approaches using diffusion models have emerged, significantly improving training stability and generation quality. Despite their success, diffusion models require numerous iterations of running neural networks to generate high-quality samples, making the increasing computational cost a potential barrier for robotics applications. To address this challenge, this paper presents R2Flow, a fast and high-fidelity generative model for LiDAR data. Our method is based on rectified flows that learn straight trajectories, simulating data generation with significantly fewer sampling steps compared to diffusion models. We also propose an efficient Transformer-based model architecture for processing the image representation of LiDAR range and reflectance measurements. Our experiments on unconditional LiDAR data generation using the KITTI-360 dataset demonstrate the effectiveness of our approach in terms of both efficiency and quality.
@@ I Introduction
LiDAR sensors provide accurate 3D point clouds of their surroundings using omnidirectional time-of-flight (ToF) ranging. The LiDAR point clouds play a crucial role in enabling autonomous mobile robots to understand their surroundings both geometrically and semantically, through techniques such as SLAM, object detection, and semantic segmentation. However, the performance of these techniques can be degraded due to incompleteness in adverse weather conditions and point density gaps between different LiDAR sensors. Restoring the degraded point clouds requires a data prior that models complex real-world patterns.
Generative modeling of LiDAR data [1, 2, 3, 4, 5, 6, 7, 8] has been studied to address this challenge, motivated by significant progress in deep generative models [9]. Deep generative models aim to build neural networks that represent the probability density distribution underlying given samples. Prior studies have demonstrated the usefulness of LiDAR generative models in tasks like sparse-to-dense completion [2, 3, 5, 4] and simulation-to-real (sim2real) domain transfer [3], which also enhance perception tasks such as semantic segmentation. Among various frameworks for the generative models, diffusion models have led to substantial improvements in the LiDAR domain [5, 4, 6, 8], offering stable training and high-quality sample generation.
Despite their success, diffusion models require significant computational costs to generate high-quality LiDAR data. In general, sampling in diffusion models is formulated as a stochastic differential equation (SDE) [10] that describes the sample trajectories from a latent distribution to the data distribution. Accurately simulating these learned trajectories requires hundreds or even thousands of discretized steps, each involving the execution of a deep neural network. Fig. 1 illustrates the trade-offs between the number of sampling steps and the sample quality for the latest methods on the LiDAR data generation [6, 4]. As shown in Fig. 1, naively reducing the number of sampling steps degrades the quality of the generated LiDAR data. This limitation poses a challenge for robotics applications, where power efficiency and computational speed are critical constraints.
LiDM [6] R2DM [4] R2Flow (ours)
Refer to captionblurred Refer to caption Refer to caption
200 steps (3.7 s) 256 steps (3.7 s) 256 steps (7.6 s)
@@
2 steps (38 ms) 2 steps (56 ms) 2 steps (62 ms)
[…]
To this end, we propose R2Flow (Range–Reflectance Flow), a novel generative model for fast and realistic LiDAR data generation. As a framework for building generative models, we employ rectified flows [12, 13], a type of conditional flow matching framework designed to train continuous normalizing flows [14, 15]. Rectified flows have been successfully applied in natural image generation tasks, such as human faces and common objects. Similar to diffusion models, the rectified flows represent the data generation as an iterative transformation using a deep neural network. However, a key distinction is that rectified flows use deterministic straight trajectories, whereas diffusion models employ stochastic curved trajectories. This makes sampling robust to the number of steps or step size, with the potential to simulate the entire trajectory in just a single step. Following relevant studies [1, 2, 3, 4, 5, 6], our R2Flow is trained on the equirectangular image representation of multimodal measurements: range and reflectance (intensity of laser reflection). We also propose a neural network architecture to generate the multimodal images based on the lightweight Vision Transformer (ViT) [16]. Among recent architectures, we verify that our approach achieves better sample quality for LiDAR data generation while reducing the computational cost and the model size. We evaluate our approach through an unconditional generation task on the KITTI-360 dataset [17]. Our approach outperforms the state-of-the-art results for both large and small numbers of steps. We summarize our contributions as follows:
@@
  • We introduce a ViT-based model architecture that balances fidelity and efficiency in LiDAR data generation.
  • We demonstrate the effectiveness of our approach through an unconditional generation task on the KITTI-360 dataset.
@@ II Related Work
Here, we briefly summarize three classes of existing LiDAR generative models using neural networks.
@@
Variational autoencoders (VAEs). VAEs [18] are trained with an autoencoder with latent representation at the bottleneck. Caccia et al. [1] initiated early work on generating LiDAR range images using a vanilla VAE. While VAEs provide stable training with the ELBO objective, they often produce blurry samples. More recently, Xiong et al. [7] employed the improved framework, a vector-quantized variational autoencoder (VQ-VAE) [19], with the voxel-based LiDAR data representation.
@@
Generative adversarial networks (GANs). GANs [20] consist of two competing neural networks, a generator and a discriminator, and have been actively applied in various domains over the last decade. Caccia et al. [1] reported the first results by training a basic GAN on range images. DUSty [2] and DUSty v2 [3] proposed architectures designed to be robust against raydrop noise (missing points caused by non-returned laser signals). Although GANs achieve better sample quality than VAEs, they suffer from unstable training and generated point clouds still deviate from real samples.
@@
Diffusion models. In recent years, diffusion models have gained significant attention for their stable training and high-quality sample generation. Diffusion models define bidirectional transitions based on a multi-step Markov process between data and latent variable spaces of the same dimensionality. Various formulations, such as score matching with Langevin dynamics (SMLD) [21, 22] and denoising diffusion probabilistic modeling (DDPM) [23, 24], have been proposed to schedule these transitions. It is also known that these formulations can be generalized as stochastic differential equations (SDEs) [10]. Several studies have applied diffusion models to LiDAR data generation. LiDARGen [5] employs SMLD, also known as a variance exploding SDE [10], to train range and reflectance images in pixel-space using a discrete-time schedule. R2DM [4] employs DDPM, also known as a variance preserving SDE [10], to also train range and reflectance images in pixel-space using a continuous-time schedule. Pixel-space diffusion models can capture fine details, but they incur high computational costs due to the iterative nature of sampling. To mitigate this issue, Ran et al. [6] proposed LiDM with architectural improvements based on the latent diffusion model (LDM) [25]. LiDM first pre-trained an autoencoder to compress the range images and then trained a discrete-time diffusion model on the lower dimensional feature space. RangeLDM [8] is a more recent work following a similar LDM approach. Nevertheless, the LDM approaches still struggle with blurriness caused by the non-iterative decoding by the autoencoder. Fig. 2 illustrates architectural comparison of the pixel-space and feature-space approaches. We prioritize the pixel precision required for range images and employ the pixel-space approach.
Refer to caption Refer to caption
Pixel-space iteration Feature-space iteration
(LiDARGen [5], R2DM [4], R2Flow) (LiDM [6], RangeLDM [8])
Figure 2: Architectural comparison of LiDAR diffusion models and ours. Our approach R2Flow is categorized into the pixel-space iteration approach.
@@
In summary, the diffusion model-based approach can generate high-fidelity samples with stable training among the generative model frameworks. However, the sampling process requires a sufficiently large number of steps because the generative process is defined by stochastic curved trajectories defined as SDEs. If these trajectories are approximated with too few steps, the generated LiDAR samples are prone to discretization errors (see Fig. 1 for an example). In this paper, we address this issue by introducing easy-to-approximate trajectories demonstrated in natural image domains [12].
@@ III Method
We employ rectified flow [12] and its extension [13] to construct generative trajectories optimized for straightness, enabling efficient sampling with only a few steps. In this section, we first introduce the procedure for building straight trajectories by rectified flows and then describe our modifications to LiDAR data generation.
@@ III-A Preliminary
Initial training. Suppose the unknown data distribution $p_{1}$ where the dataset samples $\bm{x}_{1}\sim p_{1}$ are only accessible, and Gaussian distribution $p_{0}=\mathcal{N}(0,\bm{I})$ which draws the latent variables $\bm{x}_{0}\sim p_{0}$. Both $p_{1}$ and $p_{0}$ are defined over $\mathbb{R}^{d}$. The goal is to build a transport map between the two distributions. In rectified flows, the data transformation is formulated as the following ordinary differential equation (ODE):
@@
where $\bm{x}_{t}\in\mathbb{R}^{d}$ is an intermediate state at timestep $t\in[0,1]$ and $v_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a neural network to predict the velocity fields towards $\bm{x}_{1}$. To minimize the discretization errors in ODE integration, a linear interpolation path is considered: $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$. Fig. 4(a) illustrates the example trajectory. Then we train the neural network $v_{\theta}$ by minimizing the following conditional flow matching loss $\mathcal{L}_{\rm{CFM}}$ for independently sampled pairs $(\bm{x}_{1},\bm{x}_{0})$, so that $v_{\theta}$ encourages the sample $\bm{x}_{t}$ to follow the uniform velocity $\bm{x}_{1}-\bm{x}_{0}$ as closely as possible [12].
@@
where $t\sim\mathrm{Uniform}(0,1)$. The initial model $v_{\theta}$ obtained from Eq. 2 is referred to as 1-RF in this paper.
@@
Sampling. Sampling can be performed by solving the initial value problem of the ODE described in Eq. 1, using the learned velocity estimators $v_{\theta}$. For 1-RF and 2-RF, we can choose the arbitrary number of sampling steps. In general, the larger the number of steps, the smaller the discretization error. For $k$-TD models, the number of sampling steps is fixed at $k$. Any integration solvers can be used for sampling, such as the following Euler method:
@@
Figure 3: Scene interpolation using R2Flow inversion. The both side were reconstructed from real samples via inversion. The middle four samples were generated using interpolated latent variables.
@@ III-B Data Representation
Following the existing studies [1, 2, 3, 4, 5, 6], R2Flow is trained on the equirectangular image representation of LiDAR data. We assume a LiDAR sensor that has an angular resolution of $W$ in azimuth and $H$ in elevation and measures the range and reflectance at each laser angle. Then, $HW$ sets of the range and reflectance values can be projected to a 2-channel equirectangular image $\bm{x}_{1}\in\mathbb{R}^{2\times{H}\times{W}}$ by spherical projection. Moreover, Following prior work [5, 4], we also rescale the range modality $\bm{x}_{\rm{range}}\in[0,x_{\rm{max}}]^{1\times{H}\times{W}}$ to a log-scale representation $\bm{x}_{\rm{log}}\in[0,1]^{1\times{H}\times{W}}$ as follows:
@@ III-C Velocity Estimator
Pixel-space vs. feature-space. As discussed in Sec. II, the pixel-space iteration involves high computational cost in general, which poses a barrier to the adoption of powerful backbone models such as Vision Transformers (ViTs) [27]. A common approach to this issue is to reduce the dimensionality at which iterative models operate. In the context of diffusion models, Rombach et al. [25] proposed LDM (latent diffusion model), which consists of an autoencoder (AE) pretrained to perceptually compress images and a diffusion model trained on the lower-dimensional AE features. The compression is motivated by the observation that representing imperceptible details of natural images can be relegated to AE. However, the AE-based non-iterative decoding can be problematic in LiDAR range image generation which requires accurate pixel values and their alignment to maintain geometric fidelity in point clouds. For instance, LiDM [6] proposed the LDM-based approach in the LiDAR generation task but also identified blurry patterns output by the AE decoder as a remaining issue (see Figs. 1 and 7). Therefore, in this paper, we reconsider powerful yet efficient pixel-space architectures for precise modeling, while minimizing the number of iterations by rectified flows.
Architecture design. Our model is built upon HDiT (hourglass diffusion transformer) [16], which is a ViT-based architecture proposed for pixel-space diffusion models. The key idea is to use a sliding window self-attention mechanism [26] to avoid increasing the $\mathcal{O}(n^{2})$ computation with respect to the number of tokens $n$. It thereby enables high-resolution image generation without requiring the AE-based compression [25] or staged upsampling [28], despite the pure Transformer structure. To adapt HDiT for panoramic LiDAR range and reflectance image generation, we introduce the following modifications. (i) The sliding window in the self-attention layers [26] is modified to operate in a horizontal circular pattern using a circular padding technique [5, 4]. (ii) Following the recent ViT-based architecture for LiDAR processing [29], the patch size in tokenization is changed from the default square shape to a landscape shape of $1\times 4$. The sliding windows also has the landscape shape of $3\times 9$. (iii) We use pre-defined LiDAR beam angles to condition the relative positional embeddings (RoPE [30]) in the self-attention layers, limiting the angular frequencies to harmonics. (iv) Similar to ViT [27], we apply a learnable additive bias to the tokens as an absolute positional embedding (APE); otherwise, the generated LiDAR point clouds involve random azimuth rotation.